You have collected the following information about Arguello Corporation.
•         Current earnings are 2.00 per share. Earnings will grow 20% annually in stage 1. Stage 1 lasts two years. Then, in stage 2, earnings will grow at 6% forever.
•         Dividends will be 30% of earnings in stage 1. In stage 2, because the company is growing more slowly, dividends will be 60% of earnings.
•         The required rate of return is 12%.
What is the value per share of Arguello?

A.

$24.05

B.

$25.67

C.

$28.33

D.

$29.17
Answer #1

Current Earnings = $2

E1 = $2 * (1.20) = $2.40

E2 = $2.40 * (1.20) = $2.88

Terminal value of earnings after year 2 = $2.88 * 1.06 / (0.12-0.06) = $50.88

D1 = $2.40 * 30% = $0.72

D2 = $2.88 * 30% = $0.864

Terminal value of Dividends = $50.88 * 60% = $30.528

Value per share = $0.72 * [1/(1 + 0.12)1] + $0.864 * [1/(1 + 0.12)2] + $30.528 * [1/(1 + 0.12)2]

Value per share = $25.67

Option 'B' is correct